OpenStudy (anonymous):

here is another question that is bugging me deifinite integral using definite integral

OpenStudy (anonymous):

\[\int\limits_{o}^{\pi}x \log(\sin x)= (-\pi^2\div 2)(\log 2)\]

OpenStudy (anonymous):

you have to proof that integral is equal to ((-pi^2)/2))*log 2

OpenStudy (anonymous):

ew

OpenStudy (lgbasallote):

have you tried using integration by parts?

OpenStudy (anonymous):

yeah integration by parts looks like the way to start with my guess is log(sinx) being u

OpenStudy (anonymous):

tried but not working using the properties I was able to reduce it to \[\int\limits_{0}^{\pi}\log \sin x\] I tried to solve this parts but i am stuck

OpenStudy (anonymous):

but is there anyway to make that prettier..it doesnt look so.

OpenStudy (anonymous):

@Davidjohn tell me this question has been kicking my butt only one i am not able to :(

OpenStudy (lgbasallote):

let u = log(sin x) and dv = xdx is this what you did?

OpenStudy (anonymous):

your substitution is wrong it would be du=cotxdx

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

it would be cot x...what's the problem with that hehe

OpenStudy (anonymous):

ah okay. it just takes a lot of integration by parts but things cancel out nicely. you change halfway through to the u being the x's so they eventually cancel completely